John puts R3000 on a second hand car and contracts to pay the balance in 24 monthly installments of 400 each,if interest is charged at a rate of 24% per annum payable monthly:
1.1. how much did the car cost when John purchased it?
1.2. how much interest does he pay?

Answer by math_tutor2020(3817)


L = loan amount = unknown
P = monthly payment = 400
i = monthly interest rate = 0.24/12 = 0.02
n = number of months = 24

Use those values in the monthly payment formula to solve for L
P = (L*i)/(1 - (1+i)^(-n))
400 = (L*0.02)/(1 - (1+0.02)^(-24))
400 = (0.02L)/(0.37827851206116)
0.02L = 0.37827851206116*400
0.02L = 151.311404824464
L = 151.311404824464/0.02
L = 7565.5702412232
L = 7565.57
This is the loan amount.

John made a down payment of R3000
He is loaned R7565.57 to pay the remaining cost of the car.

The price of the car is 3000+7565.57 = R10,565.57

===========================================================

John makes 24 payments of R400 each
He pays a total of 24*400 = R9600
This total consists of principal (the amount he was loaned) and interest.

Subtract off the loan amount to find the interest
9600 - 7565.57 = R2034.43 is the interest.

The down payment is NOT part of this calculation.
As far as the bank is concerned, the car costs R7565.57

===========================================================

Answers:
The price of the car is R10,565.57
The total interest paid is R2034.43